# and Qian Liu

### v1

@@ III-C.2 Medium-timescale Analog Beamformer Design
$\max_{\mathbf{F}_{\mathrm{RF}}}$ $\ \ \sum^{K}_{k=1}\widetilde{R}_{k}$ (42a)
@@
where (42a) defines the ergodic sum-rate $\widetilde{R}_{k}\triangleq\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{R_{k}\}$, evaluated over the RF domain channel samples, denoted as $\mathbf{H}_{\mathrm{RF}}\triangleq[\mathbf{h}_{\mathrm{RF},1},\mathbf{h}_{\mathrm{RF},2},\ldots,\mathbf{h}_{\mathrm{RF},K}]\in\mathbb{C}^{N_{\mathrm{t}}\times K}$ with $\mathbf{h}_{\mathrm{RF},k}$ given in (37). Notably, in this RF domain analog beamformer design problem, the power constraint is not explicitly imposed since the digital beamformer can be subsequently adjusted to satisfy it.
@@
where $\mathbf{S}_{\mathrm{RF}}\in\{0,1\}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}2^{B}}$ is the RF domain phase-shift selection matrix and $\mathbf{F}_{\mathrm{set}}$ is defined as $\mathbf{F}_{\mathrm{set}}\triangleq\mathbf{I}_{N_{\mathrm{RF}}}\otimes\mathbf{f}_{\mathrm{set}}$ with
$\mathbf{f}_{\mathrm{set}}\triangleq\frac{1}{\sqrt{N_{\mathrm{t}}}}\left[1,e^{-\jmath\frac{2\pi}{2^{B}}},\ldots,e^{-\jmath\frac{2\pi(2^{B}-1)}{2^{B}}}\right]^{T},$ (44)
containing all possible phase-shifts for a low-resolution PS. The binary matrix $\mathbf{S}_{\mathrm{RF}}$ encodes phase-shift selections, where $\mathbf{S}_{\mathrm{RF}}(n,l_{\mathrm{RFB}})=1$ indicates that the $n$-th antenna is assigned to the $l_{\mathrm{RFB}}/2^{B}$-th RF chain with the corresponding phase-shift value given by $\mathbf{F}_{\mathrm{set}}(l_{\mathrm{RFB}},l_{\mathrm{RFB}}/2^{B})$. With $\mathbf{F}_{\mathrm{set}}$ predefined, the analog beamformer design problem (42) can be equivalently formulated as
$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ \sum^{K}_{k=1}\widetilde{R}_{k}$ (45a)
@@
$\widetilde{R}_{k}=\log(1+\frac{|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}}),~{}\forall k.\$ (46)
@@
The primary challenges in solving this problem stem from the stochastic nature of the objective function (45a), which lacks a closed-form expression, and the non-smooth, non-convex Boolean constraint (45b). To address these issues, we employ a stochastic successive convex approximation (SSCA) framework [25], which iteratively refines the beamformer based on observed channel realizations. The expectation in the objective function is approximated using a quadratic surrogate function, allowing for online adaptation without requiring explicit statistical knowledge of the channel. Furthermore, the Boolean constraint is handled using a penalty-based approach, similar to the method described in Section III-B.
@@
First, we construct a recursive quadratic surrogate function for the objective function (45a), leveraging both current and past channel samples. This iterative approach enables the BS to gradually learn the statistical properties of the channel by continuously accumulating channel observations, thereby enhancing beamforming performance over time and ensuring convergence. For convenience in the following expressions, we define $g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\triangleq R_{k}$, and introduce $\mathbf{s}_{\mathrm{RF}}\triangleq\mathrm{vec}(\mathbf{S}_{\mathrm{RF}})$.
@@
At the $t_{\mathrm{M}}$-th recursion, based on the acquired channel sample $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ and the previous analog beamformer $\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$, we can substitute (45a) with a quadratic surrogate function as follows
$f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})\!=\!v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}+(\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})})^{T}(\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})})-\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\|^{2},$ (47)
@@
where $\tau>0$ is a constant ensuring strong convexity of $\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\|^{2}$. The term $v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ serves as an approximation of the objective function value $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$, which is updated recursively as
$v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}),$ (48)
@@
where $\eta^{(t_{\mathrm{M}})}\in(0,1]$ is a properly chosen sequence to satisfy convergence requirements. Similarly, the gradient term $\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ is an approximation of the partial derivative of $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$ with respect to $\mathbf{s}_{\mathrm{RF}}$, evaluated at $\mathbf{s}_{\mathrm{RF}}=\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$. This is recursively computed as
$\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}).$ (49)
@@
$\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})=\sum_{k=1}^{K}\left(\frac{\sum_{i}^{K}\mathbf{e}_{k,i}}{\mit{\Gamma_{k}}}-\frac{\sum_{i\neq k}^{K}\mathbf{e}_{k,i}}{\mit{\Gamma_{-k}}}\right),$ (50)
@@
$\mit{\Gamma_{k}}\!$ $=\!\sum_{i}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~{}\forall k,$ (51a)
$\mit{\Gamma_{-k}}\!$ $=\!\sum_{i\neq k}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~{}\forall k,$ (51b)
@@
$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})+\varrho_{2}(2\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}})}-\mathbf{1})^{T}\mathbf{s}_{\mathrm{RF}}$ (52a)
@@
$\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}=(1-\eta^{(t_{\mathrm{M}})})\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}+\eta^{(t_{\mathrm{M}})}\mathbf{\overline{S}}^{(t_{\mathrm{M}})}_{\mathrm{RF}}.$ (53)
With the updated phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$, the analog beamformer is then constructed as $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}=\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}\mathbf{F}_{\mathrm{set}}$.
@@ III-D Short-timescale Digital Beamformer Design
$\mathbf{\widetilde{F}}_{\mathrm{BB}}^{(t_{\mathrm{S}})}=(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}\left(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}+\sigma_{k}^{2}\mathbf{I}_{K}\right)^{-1}.$ (54)
Finally, to ensure compliance with the total power constraint, the digital beamformer is normalized as
@@
    • Derive the APS $\rho^{(t_{\mathrm{L}})}_{k}(\theta),\forall k$ as (19).
@@
    • Update the surrogate function (47) based on $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})},\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$.
@@
    • Update $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$ according to (53).
    • Construct the analog beamformer $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}}+1)}$ by (43).
@@ III-E.1 Summary
Based on the above derivations, the proposed tri-timescale tri-hybrid beamforming design is straightforward and summarized in Algorithm 1. After appropriately initializing the beamforming matrices $\mathbf{F}_{\mathrm{EM}}$, $\mathbf{F}_{\mathrm{RF}}$ and $\mathbf{F}_{\mathrm{BB}}$ across the three domains, the algorithm iteratively acquires the corresponding channel information and updates these matrices over the three timescales, respectively.

### v2

@@ III-C.2 Medium-timescale Analog Beamformer Design
$\max_{\mathbf{F}_{\mathrm{RF}}}$ $\ \ \sum^{K}_{k=1}\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{R_{k}\}$ (46a)
@@
where the objective function (46a) defines the ergodic sum-rate evaluated over the RF domain channel samples, denoted as $\mathbf{H}_{\mathrm{RF}}\triangleq[\mathbf{h}_{\mathrm{RF},1},\mathbf{h}_{\mathrm{RF},2},\ldots,\mathbf{h}_{\mathrm{RF},K}]\in\mathbb{C}^{N_{\mathrm{t}}\times K}$ with $\mathbf{h}_{\mathrm{RF},k}$ given in (37). Notably, in this RF domain analog beamformer design problem, the power constraint is not explicitly imposed since the digital beamformer can be subsequently adjusted to satisfy it.
@@
where $\mathbf{S}_{\mathrm{RF}}\in\{0,1\}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}2^{B}}$ is the RF domain phase-shift selection matrix and $\mathbf{F}_{\mathrm{set}}$ is defined as $\mathbf{F}_{\mathrm{set}}\triangleq\mathbf{I}_{N_{\mathrm{RF}}}\otimes\mathbf{f}_{\mathrm{set}}$ with $\mathbf{f}_{\mathrm{set}}$ containing all possible quantized phase-shifts, yielding
$\mathbf{f}_{\mathrm{set}}\triangleq\frac{1}{\sqrt{N_{\mathrm{t}}}}\left[1,e^{-\jmath\frac{2\pi}{2^{B}}},\ldots,e^{-\jmath\frac{2\pi(2^{B}-1)}{2^{B}}}\right]^{T}.$ (48)
The binary matrix $\mathbf{S}_{\mathrm{RF}}$ encodes phase-shift selections, where $\mathbf{S}_{\mathrm{RF}}(n,l_{\mathrm{RFB}})=1$ indicates that the $n$-th antenna is assigned to the $l_{\mathrm{RFB}}/2^{B}$-th RF chain with the corresponding phase-shift value given by $\mathbf{F}_{\mathrm{set}}(l_{\mathrm{RFB}},l_{\mathrm{RFB}}/2^{B})$. With $\mathbf{F}_{\mathrm{set}}$ predefined, the analog beamformer design problem (46) can be equivalently formulated as
$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ \sum^{K}_{k=1}\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{R_{k}\}$ (49a)
@@
$R_{k}\!\!=\!\log_{2}\!\left(\!1\!\!+\!\frac{|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}\!|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},j}|^{2}\!\!+\!\!\sigma_{k}^{2}}\!\right)\!,\forall k.\!\!\!\!\\$ (50)
@@
The primary challenges in solving this problem stem from the stochastic nature of the objective function (49a), which lacks a closed-form expression, and the non-smooth, non-convex Boolean constraint (49b). To address these issues, we employ a stochastic successive convex approximation (SSCA) framework [28], which recursively refines the beamformer based on observed channel realizations. The expectation in the objective function is approximated using a quadratic surrogate function, allowing for online adaptation without requiring explicit statistical knowledge of the channel. Furthermore, the Boolean constraint is handled using a penalty-based approach, similar to the method described in Section III-B.
@@
First, we construct a recursive quadratic surrogate function for the objective function (49a), leveraging both current and past channel samples. This recursive approach enables the BS to gradually learn the statistical properties of the channel by continuously accumulating channel observations, thereby enhancing beamforming performance over time and ensuring convergence. For convenience in the following expressions, we define $g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\triangleq R_{k}$, and introduce $\mathbf{s}_{\mathrm{RF}}\triangleq\mathrm{vec}(\mathbf{S}_{\mathrm{RF}})$.
@@
At the $t_{\mathrm{M}}$-th recursion, based on the acquired channel sample $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ and the previous RF domain phase-shift selection matrix $\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}$, we can substitute (49a) with a quadratic surrogate function as follows
$f^{(t_{\mathrm{M}})}\!(\mathbf{s}_{\mathrm{RF}}\!)\!=\!v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}+(\!\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}\!)^{T}\!(\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}\!-\!1)})-\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}\!-\!1)}\!\|^{2},$ (51)
@@
where $\tau>0$ is a constant ensuring strong convexity of $\tau\|\mathbf{s}_{\mathrm{RF}}-\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}\|^{2}$. The term $v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ serves as an approximation of the objective function value $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$, which is updated recursively as
$v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})v_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}),$ (52)
@@
where $\eta^{(t_{\mathrm{M}})}$ is a properly chosen sequence to satisfy convergence requirements [28]. Similarly, the gradient term $\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}$ is an approximation of the partial derivative of $\mathbb{E}_{\mathbf{H}_{\mathrm{RF}}}\{g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})\}$ with respect to $\mathbf{s}_{\mathrm{RF}}$, evaluated at $\mathbf{s}_{\mathrm{RF}}=\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}$. This is recursively computed as
$\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}})}\!\!=\!\!(1\!-\!\eta^{(t_{\mathrm{M}})})\mathbf{v}_{\mathbf{s}_{\mathrm{RF}}}^{(t_{\mathrm{M}}\!-\!1)}\!+\!\eta^{(t_{\mathrm{M}})}\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}\!-\!1)};\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})}).$ (53)
@@
$\nabla_{\mathbf{s}_{\mathrm{RF}}}g_{0}(\mathbf{s}_{\mathrm{RF}};\mathbf{H}_{\mathrm{RF}})=\sum_{k=1}^{K}\left(\frac{\sum_{i}^{K}\mathbf{e}_{k,i}}{{\Gamma_{k}}}-\frac{\sum_{i\neq k}^{K}\mathbf{e}_{k,i}}{{\Gamma_{-k}}}\right),$ (54)
@@
${\Gamma_{k}}\!$ $=\!\sum_{i}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~\forall k,$ (55a)
${\Gamma_{-k}}\!$ $=\!\sum_{i\neq k}^{K}\left|\mathbf{h}_{\mathrm{RF},k}^{H}\mathbf{S}_{\mathrm{RF}}\mathbf{F}_{\mathrm{set}}\mathbf{f}_{\mathrm{BB},i}\right|^{2}+\sigma_{k}^{2},~\forall k,$ (55b)
@@
$\max_{\mathbf{S}_{\mathrm{RF}}}$ $\ \ \ f^{(t_{\mathrm{M}})}(\mathbf{s}_{\mathrm{RF}})+\varrho_{2}(2\mathbf{s}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}-\mathbf{1})^{T}\mathbf{s}_{\mathrm{RF}}$ (56a)
@@
$\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}=(1-\eta^{(t_{\mathrm{M}})})\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}+\eta^{(t_{\mathrm{M}})}\mathbf{\overline{S}}^{(t_{\mathrm{M}})}_{\mathrm{RF}}$. (57)
With the updated phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$, the analog beamformer is then constructed as $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}})}=\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}\mathbf{F}_{\mathrm{set}}$.
@@ III-D Short-timescale Digital Beamformer Design
$\mathbf{\widetilde{F}}_{\mathrm{BB}}^{(t_{\mathrm{S}})}=(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}\left(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}(\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})})^{H}+\bm{\Lambda}\right)^{-1},$ (58)
where $\bm{\Lambda}\triangleq\mathrm{diag}\{\sigma_{1}^{2},\sigma_{2}^{2},\ldots,\sigma_{K}^{2}\}$. Finally, to ensure compliance with the total power constraint, the digital beamformer is normalized as
@@
    • Derive the APS $\rho^{(t_{\mathrm{L}})}_{k}(\vartheta_{m}),\forall k$ as (20).
@@
    • Update the surrogate function (51) based on $\mathbf{H}_{\mathrm{RF}}^{(t_{\mathrm{M}})},\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}}-1)}$.
@@
    • Update $\mathbf{S}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ according to (57).
    • Construct the analog beamformer $\mathbf{F}_{\mathrm{RF}}^{(t_{\mathrm{M}})}$ by (47).
@@ III-E.1 Summary
Based on the above derivations, the proposed tri-timescale tri-hybrid beamforming design is straightforward and summarized in Algorithm 1. After appropriately initializing the beamforming matrices $\mathbf{F}_{\mathrm{EM}}$, $\mathbf{F}_{\mathrm{RF}}$ and $\mathbf{F}_{\mathrm{BB}}$ across the three domains, the algorithm alternately acquires the corresponding channel information and updates these matrices over the three timescales, respectively.